# Deep Learning for 3D Human Pose Estimation and Mesh Recovery: A Survey

### v1

@@ 5.1.1 Naked human body recovery
Developing Various Representations. Stable and effective feature representation is crucial for enhancing the capabilities of deep learning algorithms in human mesh recovery, as it enables the efficient extraction of meaningful patterns from complex input data. Based on the assumption that image texture remains constant across frames, TexturePose pavlakos2019texturepose leverages the appearance constancy of the body across different frames. It measures the texture map for each frame using a texture consistency loss. DenseRaC xu2019denserac generates a pixel-to-surface correspondence map to optimize the estimation of parameterized human pose and shape. Zhang et al. zhang2020learning established a connection between 2D pixels and 3D vertices by using dense correspondences of body parts, effectively addressing related issues. Their proposed DaNet model concentrates on learning the 2D-to-3D mapping, while the PartDrop strategy ensures that the model focuses more on complementary body parts and adjacent positional features. To address the lack of dense correspondence between image features and the 3D mesh, DecoMR zeng20203d recovers the human mesh by establishing a pixel-to-surface dense correspondence map in the UV space. Zhang et al. zhang2020object designed a two-branch network that utilizes a partial UV map to represent the human body when occluded by objects, effectively converting this map into an estimation of the 3D human shape. Sun et al. sun2021monocular proposed a body-center-guided representation method that predicts both the body center heatmap and the mesh parameter map. This approach describes the 3D body mesh at the pixel level, enabling one-stage multi-person 3D mesh regression. 3DCrowdNet choi2022learning employs a joint-based regressor to isolate target features from others for recovering multi-person meshes in crowded scenes, which utilizes a crowded scene-robust image feature heatmap instead of the full feature map within a bounding box. Dual-stream Spatio-temporal Transformer (DSTformer) zhu2023motionbert extracts long-range spatio-temporal relationships among skeletal joints to effectively capture unified human motion representations from large-scale and heterogeneous video data for human-centric tasks, such as human mesh recovery.
@@
Utilizing Structural Information. The structural information of the body acts as unique prior knowledge in human mesh recovery, enhancing the understanding of body relations. Furthermore, introducing additional physical constraints, which describe the interrelationships between various body structures, can significantly improve the performance. Holopose guler2019holopose employs a part-based multi-task regression network for 2D, 3D, and dense pose to estimate 3D human surfaces. Sun et al. sun2019human introduced an end-to-end method for human mesh recovery from single images and monocular videos. This method employs skeleton disentangling to reduce the complexity of decoupling and incorporates temporal coherence, efficiently capturing both short and long-term temporal cues. Hybrid Inverse Kinematics (HybrIK) li2021hybrik calculates the swing rotation from 3D joints and employs a network to predict the twist rotation through the twist-and-swing decomposition. The method can tackle non-linearity challenges and misalignment between images and models in mesh estimation from 3D poses. In their later work, Li et al. li2023niki designed NIKI, a model capable of learning from both the forward and inverse processes using invertible networks. To address the non-linear mapping and drifting joint position issues, Lee et al. lee2021uncertainty introduced an uncertainty-aware method for human mesh recovery, leveraging information from 2D poses to address the inherent ambiguities in 2D. Sengupta et al. sengupta2021hierarchical presented a probabilistic approach to circumvent the ill-posed problem, which integrates the kinematic tree structure of the human body with a Gaussian distribution over SMPL parameters. This method then predicts the hierarchical matrix-fisher distribution of 3D joint rotation matrices. SGRE wang20233d estimates the global rotation matrix of joints directly to avoid error accumulation along the kinematic chains in human mesh recovery.
@@
Choosing Appropriate Learning Strategies. SPIN kolotouros2019learning incorporates an initial estimate optimization routine into the training loop by the self-improving neural network, which can fit the body mesh estimate to 2D joints. Zanfir et al. zanfir2020weakly developed a method integrating kinematic latent normalizing flow representations and dynamical models with structured, differentiable, semantic body part alignment loss functions aimed at enhancing semi-supervised and self-supervised 3D human pose and shape estimation. Jiang et al. jiang2020coherent introduced two novel loss functions for multi-person mesh recovery from single images: a distance field-based collision loss penalizing interpenetration between constructed figures and a depth ordering-aware loss addressing occlusions and promoting accurate depth ordering of targets. Madadi et al. madadi2021deep presented an unsupervised denoising autoencoder network to recover invisible landmarks using sparse motion capture data effectively. To tackle the challenges of pose failure and shape ambiguity in the unsupervised human mesh recovery task, Yu et al. yu2021skeleton2mesh devised a strategy that decouples the task into unsupervised 3D pose estimation and leverages kinematic prior knowledge. Bilevel Online Adaptation (BOA) guan2021bilevel employs bilevel optimization to reconcile conflicts between 2D and temporal constraints in out-of-domain streaming videos human mesh recovery. In their later work, Dynamic Bilevel Online Adaptation (DBOA) guan2022out integrates temporal constraints to compensate for the absence of 3D annotations. Huang et al. huang2022pose2uv developed Pose2UV, a single-shot human mesh recovery method capable of extracting target features under occlusions using a deep UV prior. JOTRli2023jotr fuses 2D and 3D features and incorporates supervision for the 3D feature through a Transformer-based contrastive learning framework. ReFit wang2023refit reprojects keypoints and refines the human model via a feedback-update loop mechanism. You et al. you2023co introduced a co-evolution method for human mesh recovery that utilizes 3D pose as an intermediary. This method divides the process into two distinct stages: initially, it estimates the 3D human pose from video, and subsequently, it regresses mesh vertices based on the estimated 3D pose, combined with temporal image features. To bridge the gap between training and test data, CycleAdapt nam2023cyclic proposed a domain adaptation method including a mesh reconstruction network and a motion denoising network enabling more effective adaptation.
@@ 5.1.2 Detailed human body recovery
Model-based human mesh recovery has moderately satisfactory results but still lacks detailed body representations; expansions on the naked parametric model now allow for parameterized depictions of various body parts and details, including clothing jiang2020bcnet ; jiang2022h4d , hands romero2022embodied , face li2017learning , and the entire body pavlakos2019expressive , as discussed in Section 2.2.
@@
With Clothes. Alldieck et al. alldieck2019learning estimated the parameters of the SMPL model, including clothing and hair, from 1 to 8 frames of a monocular video. Bhatnagar et al. bhatnagar2019multi developed the Multi-Garment Network (MGN) to reconstruct body shape and clothing layers on top of the SMPL model. Tex2Shape alldieck2019tex2shape converts the human body mesh regression problem into an image-to-image estimation task. Specifically, it predicts a partial texture map of the visible region and then reconstructs the body shape, adding details to visible and occluded parts. BCNet jiang2020bcnet features a layered garment representation atop the SMPL model and innovatively decouples the skinning weight of the garment from the body mesh. Jiang et al. jiang2022h4d introduced a method that employs a temporal span, SMPL parameters of shape and initial pose, and latent codes encoding motion and auxiliary information. This approach facilitates the recovery of detailed body shapes, including visible and occluded parts, by utilizing a texture map of the visible region.
@@
With Hands. Forte et al. forte2023reconstructing proposed SGNify, a model that captures hand pose, facial expression, and body movement from sign language videos. It employs linguistic priors and constraints on 3D hand pose to effectively address the ambiguities in isolated signs. Additionally, the relationship between Two-Hands zhang2021interacting , and Hand-Object chen2021joint effectively reconstructs the hand’s details.
@@
Whole Body. To address the inconsistency between 3D human mesh and 3D scenes, Hassan et al. hassan2019resolving introduced a human whole body and scenes recovery method named Proximal Relationships with Object eXclusion (PROX). EXpressive POse and Shape rEgression (ExPose) framework choutas2020monocular employs a body-driven attention mechanism and adopts a regression approach for holistic expressive body reconstruction to mitigate local optima issues in optimization-based methods. FrankMocap rong2021frankmocap operates by independently running 3D mesh recovery regression for face, hands, and body and subsequently combining the outputs through an integration module. PIXIE feng2021collaborative integrates independent estimates from the body, face, and hands using the shared shape space of SMPL-X across all body parts. Moon et al. moon2022accurate developed an end-to-end framework for whole-body human mesh recovery named Hand4Whole, which employs joint features for 3D joint rotations to enhance the accuracy of 3D hand predictions. Zhang et al. zhang2023pymaf enhanced the PyMAF framework zhang2021pymaf , developing PyMAF-X for the detailed reconstruction of full-body models. This advancement aims to resolve the misalignment issues in regression-based, one-stage human mesh recovery methods by employing a feature pyramid approach and refining the mesh-image alignment parameters. OSX lin2023one employs a simple yet effective component-aware transformer that includes a global body encoder and a local face/hand decoder instead of separate networks for each part. Li et al. li2023hybrik extended the HybrIK li2021hybrik framework and proposed HybrIK-X, a one-stage model based on a hybrid analytical-neural inverse kinematics framework to recover comprehensive whole-body meshes with details.
@@ 5.2 Template-free human body recovery
Regression-based human mesh recovery bypasses the limitations and biases inherent in template-based models and directly outputs the template-free 3D models of the human body. This approach allows for a more dynamic and flexible generation of models, which can capture a wider variety of human body shapes and postures. FACSIMILE (FAX) utilizes an image-translation network to recover geometry at the original image resolution directly, bypassing the need for indirect output through representations. Jinka et al. jinka2020peeledhuman proposed a robust shape representation specifically for scenes with self-occlusions, where the method encodes the human body using peeled RGB and depth maps, significantly enhancing accuracy and efficiency during both training and inference. Neural Surface Fields (NSF) xue2023nsf models a continuous and flexible displacement field on the base surface for 3D clothed human mesh recovery from monocular depth. This approach adapts to base surfaces with varying resolutions and topologies without retraining during inference. Zhang et al. zhang2023global introduced the Global-correlated 3D-decoupling Transformer for clothed Avatar reconstruction (GTA), a model that employs an encoder to capture globally-correlated image features and a decoder to decouple tri-plane features using cross-attention. Optimization-based differentiable rendering integrates the rendering process into the optimization method by minimizing the rendering error. AG3D dong2023ag3d captures the body’s and loose clothing’s shape and deformation by adopting a holistic 3D generator and integrating geometric cues in the form of predicted 2D normal maps. Gärtner et al. gartner2022differentiable designed DiffPhy, a differentiable physics-based model that incorporates a physically plausible body representation with anatomical joint limits, significantly enhancing the performance and robustness of human body recovery.
@@
Implicit representations do not directly represent an object’s geometric data, such as vertices or meshes, but rather define whether a point in space belongs to the object through a function. The advantage of implicit representations lies in their ability to compactly and flexibly represent complex shapes, including those with intricate topologies or discontinuities. Thus, they have achieved impressive outcomes in human body reconstruction. Saito et al. saito2019pifu proposed the Pixel-aligned Implicit Function (PIFu), an implicit representation that aligns pixels of 2D images with a 3D object, and designed an end-to-end deep learning framework for inferring both 3D surface and texture from a single image. PIFu was the first to apply implicit representation in human mesh recovery, enabling the reconstruction of geometric details of the human body in clothing. Subsequently, PIFuHD saito2020pifuhd extended the work of PIFu to 4K resolution images, further enhancing detail. Huang et al. huang2020arch designed the Animatable Reconstruction of Clothed Humans (ARCH) method by creating a semantic space and a semantic deformation field. He et al. he2021arch++ introduced ARCH++, a co-supervising framework with cross-space consistency, which jointly estimates occupancy in both the posed and canonical spaces. They transform the human mesh recovery problem with implicit representation from pose space to canonical space for processing. However, this approach’s limitation is that it heavily relies on accurate pose estimation, and the clothing expression based on skinning weights still lacks sufficient naturalness in detail. GeoPIFu he2020geo learns latent voxel features using a structure-aware 3D U-Net incorporating geometric priors. To tackle the ill-posed problem in representation learning when views are incredibly sparse, Peng et al. peng2021neural developed a novel human body representation. This representation assumes that the neural representations learned at different frames share latent codes anchored to a deformable mesh. Clothed Avatar Reconstruction (CAR) method liao2023high employs a learning-based implicit model to initially form the general shape in canonical space, followed by refining surface details through predicting non-rigid optimization deformation in the posed space. 3D-aware Neural Body Fitting (3DNBF) zhang20233d addresses the challenges of occlusion and 2D-3D ambiguity by a volumetric human representation using Gaussian ellipsoidal kernels. As illustrated in Fig.7(a), Neural Radiance Fields mildenhall2021nerf is based on the principle of implicit representation, using neural networks to learn the continuous volumetric density and color distribution of a scene, allowing for generating a high-quality 3D model from arbitrary viewpoints. Gao et al. gao2022mps utilized a specialized representation that combines a canonical Neural Radiance Field (NeRF) with a volume deformation scheme, enabling the recovery of novel views and poses for a person not seen during training. Mu et al. mu2023actorsnerf introduce ActorsNeRF, a NeRF-based human representation for human mesh recovery from a few monocular images by encoding the category-level prior through parameter sharing with a 2-level canonical space. Diffusion models are based on a series of diffusion processes, transforming original data by adding random noise and then gradually removing this noise to generate new data through a reverse process. Human Mesh Diffusion (HMDiff) foo2023distribution treats mesh recovery as a reverse diffusion process, incorporates input-specific distribution information into the diffusion process, and introduces prior knowledge to simplify the task.
@@
Methods based on implicit representations can recover free-form geometric shapes, but they may have excessive dependence on accurate poses or generate unsatisfactory shapes for novel poses or clothing. To increase robustness in these scenarios, existing work employs explicit body models to constrain mesh reconstruction in implicit methods. Researchers desire an approach that combines implicit and explicit methods based on a coarse body mesh’s geometry to refine and produce detailed human models meticulously. Hierarchical Mesh Deformation (HMD) zhu2019detailed uses constraints from body joints, silhouettes, and per-pixel shading information to combine the robustness of a parametric model with the flexibility of free-form 3D deformation. Bhatnagar et al. bhatnagar2020combining introduced an Implicit Part Network (IP-Net) to predict the detailed human surface, including the 3D surface of the dressed person, the inner body surface, and the parametric body model. Parametric Model-Conditioned Implicit Representation (PaMIR) zheng2021pamir combines a parametric body model with a free-form deep implicit function to enhance generalization through regularization. Zhu et al. zhu2021detailed utilized a ’project-predict-deform’ strategy that refines the SMPL model generated by human recovery methods using supervision from joints, silhouettes, and shading information. ICON xiu2022icon generates a stable, coarse human mesh using the SMPL model, then renders the front and back body normals, which provide rich texture details and combine with the original image. Finally, the body’s normal clothed normal, and Signed Distance Function (SDF) information are fed into an implicit MLP to obtain the final result. Subsequently, Xiu et al. proposed ECON xiu2023econ , which integrates normal estimation, normal integration, and shape completion by using SMPL-X depth as a soft geometric constraint within the optimization equation of Normal Integration. It endeavors to maintain coherence with nearby surfaces during the integration of normals. Feng et al. feng2023learning presented Disentangled Avatars (DELTA), a model representing humans with hybrid explicit-implicit 3D representations. GETAvatar zhang2023getavatar directly produces explicit textured 3D human meshes. GETAvatar creates an articulated 3D human representation with explicit surface modeling, enriches it with realistic surface details derived from 2D normal maps of 3D scan data, and utilizes a rasterization-based renderer for surface rendering. Diffusion Inpainting of Neural Avatars (DINAR) svitov2023dinar combines neural textures with the SMPL-X body model and employs a latent diffusion model to recover textures of both seen and unseen regions, subsequently integrating these onto the base SMPL-X mesh. TransHuman pan2023transhuman employs transformers to project the SMPL model into a canonical space and associates each output token with a deformable radiance field. This field encodes the query point in the observation space and is further utilized to integrate fine-grained information from reference images. Gaussian Splatting kerbl20233d is a technique for processing and rendering point clouds, using a Gaussian function to create an influence range for each point, resulting in a smoother and more natural representation in two-dimensional images, as shown in Fig.7(b). It has achieved good results in SLAM (Simultaneous Localization and Mapping) yan2023gs , generative human modeling liu2023humangaussian , dynamic scene reconstruction wu20234d , and multimodal generation chen2023text . Reconstruction based on Gaussian Splatting requires an initial point cloud input, which poses a challenge for human mesh recovery from image inputs. However, explicit models can provide initial points, serving as a stepping stone to make Gaussian splatting human mesh recovery from RGB inputs feasible. Animatable 3D Gaussian liu2023animatable learns human avatars from input images and poses by extending 3D Gaussian to dynamic human scenes. In the proposed framework, they model a set of skinned 3D Gaussian and a corresponding skeleton in canonical space and deform 3D Gaussian to posed space according to the input poses.

### v2

@@ 5.1.1 Naked human body recovery
Developing Various Representations. Stable and effective feature representation is crucial for enhancing the capabilities of deep learning algorithms in human mesh recovery, as it enables the efficient extraction of meaningful patterns from complex input data. Based on the assumption that image texture remains constant across frames, TexturePose [150] leverages the appearance constancy of the body across different frames. It measures the texture map for each frame using a texture consistency loss. DenseRaC [216] generates a pixel-to-surface correspondence map to optimize the estimation of parameterized human pose and shape. Zhang et al. [151] established a connection between 2D pixels and 3D vertices by using dense correspondences of body parts, effectively addressing related issues. Their proposed DaNet model concentrates on learning the 2D-to-3D mapping, while the PartDrop strategy ensures that the model focuses more on complementary body parts and adjacent positional features. To address the lack of dense correspondence between image features and the 3D mesh, DecoMR [152] recovers the human mesh by establishing a pixel-to-surface dense correspondence map in the UV space. Zhang et al. [153] designed a two-branch network that utilizes a partial UV map to represent the human body when occluded by objects, effectively converting this map into an estimation of the 3D human shape. Sun et al. [154] proposed a body-center-guided representation method that predicts both the body center heatmap and the mesh parameter map. This approach describes the 3D body mesh at the pixel level, enabling one-stage multi-person 3D mesh regression. 3DCrowdNet [155] employs a joint-based regressor to isolate target features from others for recovering multi-person meshes in crowded scenes, which utilizes a crowded scene-robust image feature heatmap instead of the full feature map within a bounding box. Dual-stream Spatio-temporal Transformer (DSTformer) [156] extracts long-range spatio-temporal relationships among skeletal joints to effectively capture unified human motion representations from large-scale and heterogeneous video data for human-centric tasks, such as human mesh recovery.
@@
Utilizing Structural Information. The structural information of the body acts as unique prior knowledge in human mesh recovery, enhancing the understanding of body relations. Furthermore, introducing additional physical constraints, which describe the interrelationships between various body structures, can significantly improve the performance. Holopose [157] employs a part-based multi-task regression network for 2D, 3D, and dense pose to estimate 3D human surfaces. Sun et al. [158] introduced an end-to-end method for human mesh recovery from single images and monocular videos. This method employs skeleton disentangling to reduce the complexity of decoupling and incorporates temporal coherence, efficiently capturing both short and long-term temporal cues. Hybrid Inverse Kinematics (HybrIK) [159] calculates the swing rotation from 3D joints and employs a network to predict the twist rotation through the twist-and-swing decomposition. The method can tackle non-linearity challenges and misalignment between images and models in mesh estimation from 3D poses. In their later work, Li et al. [160] designed NIKI, a model capable of learning from both the forward and inverse processes using invertible networks. To address the non-linear mapping and drifting joint position issues, Lee et al. [161] introduced an uncertainty-aware method for human mesh recovery, leveraging information from 2D poses to address the inherent ambiguities in 2D. Sengupta et al. [162] presented a probabilistic approach to circumvent the ill-posed problem, which integrates the kinematic tree structure of the human body with a Gaussian distribution over SMPL parameters. This method then predicts the hierarchical matrix-fisher distribution of 3D joint rotation matrices. SGRE [163] estimates the global rotation matrix of joints directly to avoid error accumulation along the kinematic chains in human mesh recovery.
@@
Choosing Appropriate Learning Strategies. SPIN [164] incorporates an initial estimate optimization routine into the training loop by the self-improving neural network, which can fit the body mesh estimate to 2D joints. Zanfir et al. [48] developed a method integrating kinematic latent normalizing flow representations and dynamical models with structured, differentiable, semantic body part alignment loss functions aimed at enhancing semi-supervised and self-supervised 3D human pose and shape estimation. Jiang et al. [166] introduced two novel loss functions for multi-person mesh recovery from single images: a distance field-based collision loss penalizing interpenetration between constructed figures and a depth ordering-aware loss addressing occlusions and promoting accurate depth ordering of targets. Madadi et al. [167] presented an unsupervised denoising autoencoder network to recover invisible landmarks using sparse motion capture data effectively. To tackle the challenges of pose failure and shape ambiguity in the unsupervised human mesh recovery task, Yu et al. [50] devised a strategy that decouples the task into unsupervised 3D pose estimation and leverages kinematic prior knowledge. Bilevel Online Adaptation (BOA) [217] employs bilevel optimization to reconcile conflicts between 2D and temporal constraints in out-of-domain streaming videos human mesh recovery. In their later work, Dynamic Bilevel Online Adaptation (DBOA) [168] integrates temporal constraints to compensate for the absence of 3D annotations. Huang et al. [169] developed Pose2UV, a single-shot human mesh recovery method capable of extracting target features under occlusions using a deep UV prior. JOTR[170] fuses 2D and 3D features and incorporates supervision for the 3D feature through a Transformer-based contrastive learning framework. ReFit [165] reprojects keypoints and refines the human model via a feedback-update loop mechanism. You et al. [5] introduced a co-evolution method for human mesh recovery that utilizes 3D pose as an intermediary. This method divides the process into two distinct stages: initially, it estimates the 3D human pose from video, and subsequently, it regresses mesh vertices based on the estimated 3D pose, combined with temporal image features. To bridge the gap between training and test data, CycleAdapt [171] proposed a domain adaptation method including a mesh reconstruction network and a motion denoising network enabling more effective adaptation.
@@ 5.1.2 Detailed human body recovery
Model-based human mesh recovery has moderately satisfactory results but still lacks detailed body representations; expansions on the naked parametric model now allow for parameterized depictions of various body parts and details, including clothing [175, 39], hands [36], face [37], and the entire body [38], as discussed in Section 2.2.
@@
With Clothes. Alldieck et al. [172] estimated the parameters of the SMPL model, including clothing and hair, from 1 to 8 frames of a monocular video. Bhatnagar et al. [173] developed the Multi-Garment Network (MGN) to reconstruct body shape and clothing layers on top of the SMPL model. Tex2Shape [174] converts the human body mesh regression problem into an image-to-image estimation task. Specifically, it predicts a partial texture map of the visible region and then reconstructs the body shape, adding details to visible and occluded parts. BCNet [175] features a layered garment representation atop the SMPL model and innovatively decouples the skinning weight of the garment from the body mesh. Jiang et al. [39] introduced a method that employs a temporal span, SMPL parameters of shape and initial pose, and latent codes encoding motion and auxiliary information. This approach facilitates the recovery of detailed body shapes, including visible and occluded parts, by utilizing a texture map of the visible region.
@@
With Hands. Forte et al. [176] proposed SGNify, a model that captures hand pose, facial expression, and body movement from sign language videos. It employs linguistic priors and constraints on 3D hand pose to effectively address the ambiguities in isolated signs. Additionally, the relationship between Two-Hands [177], and Hand-Object [178] effectively reconstructs the hand’s details.
@@
Whole Body. To address the inconsistency between 3D human mesh and 3D scenes, Hassan et al. [179] introduced a human whole body and scenes recovery method named Proximal Relationships with Object eXclusion (PROX). EXpressive POse and Shape rEgression (ExPose) framework [180] employs a body-driven attention mechanism and adopts a regression approach for holistic expressive body reconstruction to mitigate local optima issues in optimization-based methods. FrankMocap [181] operates by independently running 3D mesh recovery regression for face, hands, and body and subsequently combining the outputs through an integration module. PIXIE [182] integrates independent estimates from the body, face, and hands using the shared shape space of SMPL-X across all body parts. Moon et al. [183] developed an end-to-end framework for whole-body human mesh recovery named Hand4Whole, which employs joint features for 3D joint rotations to enhance the accuracy of 3D hand predictions. Zhang et al. [184] enhanced the PyMAF framework [218], developing PyMAF-X for the detailed reconstruction of full-body models. This advancement aims to resolve the misalignment issues in regression-based, one-stage human mesh recovery methods by employing a feature pyramid approach and refining the mesh-image alignment parameters. OSX [185] employs a simple yet effective component-aware transformer that includes a global body encoder and a local face/hand decoder instead of separate networks for each part. Li et al. [186] extended the HybrIK [159] framework and proposed HybrIK-X, a one-stage model based on a hybrid analytical-neural inverse kinematics framework to recover comprehensive whole-body meshes with details.
@@ 5.2 Template-free human body recovery
Regression-based Methods. Regression-based human mesh recovery bypasses the limitations and biases inherent in template-based models and directly outputs the template-free 3D models of the human body. This approach allows for a more dynamic and flexible generation of models, which can capture a wider variety of human body shapes and postures. FACSIMILE (FAX) utilizes an image-translation network to recover geometry at the original image resolution directly, bypassing the need for indirect output through representations. Jinka et al. [188] proposed a robust shape representation specifically for scenes with self-occlusions, where the method encodes the human body using peeled RGB and depth maps, significantly enhancing accuracy and efficiency during both training and inference. Neural Surface Fields (NSF) [190] models a continuous and flexible displacement field on the base surface for 3D clothed human mesh recovery from monocular depth. This approach adapts to base surfaces with varying resolutions and topologies without retraining during inference. Zhang et al. [189] introduced the Global-correlated 3D-decoupling Transformer for clothed Avatar reconstruction (GTA), a model that employs an encoder to capture globally-correlated image features and a decoder to decouple tri-plane features using cross-attention.
Optimization-based Methods. Optimization-based differentiable rendering integrates the rendering process into the optimization method by minimizing the rendering error. AG3D [192] captures the body’s and loose clothing’s shape and deformation by adopting a holistic 3D generator and integrating geometric cues in the form of predicted 2D normal maps. Gärtner et al. [191] designed DiffPhy, a differentiable physics-based model that incorporates a physically plausible body representation with anatomical joint limits, significantly enhancing the performance and robustness of human body recovery.
@@
Implicit Representation Methods. Implicit representations do not directly represent an object’s geometric data, such as vertices or meshes, but rather define whether a point in space belongs to the object through a function. The advantage of implicit representations lies in their ability to compactly and flexibly represent complex shapes, including those with intricate topologies or discontinuities. Thus, they have achieved impressive outcomes in human body reconstruction. Saito et al. [193] proposed the Pixel-aligned Implicit Function (PIFu), an implicit representation that aligns pixels of 2D images with a 3D object, and designed an end-to-end deep learning framework for inferring both 3D surface and texture from a single image. PIFu was the first to apply implicit representation in human mesh recovery, enabling the reconstruction of geometric details of the human body in clothing. Subsequently, PIFuHD [194] extended the work of PIFu to 4K resolution images, further enhancing detail. Huang et al. [195] designed the Animatable Reconstruction of Clothed Humans (ARCH) method by creating a semantic space and a semantic deformation field. He et al. [196] introduced ARCH++, a co-supervising framework with cross-space consistency, which jointly estimates occupancy in both the posed and canonical spaces. They transform the human mesh recovery problem with implicit representation from pose space to canonical space for processing. However, this approach’s limitation is that it heavily relies on accurate pose estimation, and the clothing expression based on skinning weights still lacks sufficient naturalness in detail. GeoPIFu [198] learns latent voxel features using a structure-aware 3D U-Net incorporating geometric priors. To tackle the ill-posed problem in representation learning when views are incredibly sparse, Peng et al. [199] developed a novel human body representation. This representation assumes that the neural representations learned at different frames share latent codes anchored to a deformable mesh. Clothed Avatar Reconstruction (CAR) method [197] employs a learning-based implicit model to initially form the general shape in canonical space, followed by refining surface details through predicting non-rigid optimization deformation in the posed space. 3D-aware Neural Body Fitting (3DNBF) [200] addresses the challenges of occlusion and 2D-3D ambiguity by a volumetric human representation using Gaussian ellipsoidal kernels.
Neural Radiance Fields. As illustrated in Fig.8(a), Neural Radiance Fields [219] is based on the principle of implicit representation, using neural networks to learn the continuous volumetric density and color distribution of a scene, allowing for generating a high-quality 3D model from arbitrary viewpoints. Gao et al. [201] utilized a specialized representation that combines a canonical Neural Radiance Field (NeRF) with a volume deformation scheme, enabling the recovery of novel views and poses for a person not seen during training. Mu et al. [51] introduce ActorsNeRF, a NeRF-based human representation for human mesh recovery from a few monocular images by encoding the category-level prior through parameter sharing with a 2-level canonical space.
Diffusion models. Diffusion models are based on a series of diffusion processes, transforming original data by adding random noise and then gradually removing this noise to generate new data through a reverse process. Human Mesh Diffusion (HMDiff) [202] treats mesh recovery as a reverse diffusion process, incorporates input-specific distribution information into the diffusion process, and introduces prior knowledge to simplify the task.
@@
Implicit Representation with Explicit Parameter. Methods based on implicit representations can recover free-form geometric shapes, but they may have excessive dependence on accurate poses or generate unsatisfactory shapes for novel poses or clothing. To increase robustness in these scenarios, existing work employs explicit body models to constrain mesh reconstruction in implicit methods. Researchers desire an approach that combines implicit and explicit methods based on a coarse body mesh’s geometry to refine and produce detailed human models meticulously. Hierarchical Mesh Deformation (HMD) [203] uses constraints from body joints, silhouettes, and per-pixel shading information to combine the robustness of a parametric model with the flexibility of free-form 3D deformation. Bhatnagar et al. [204] introduced an Implicit Part Network (IP-Net) to predict the detailed human surface, including the 3D surface of the dressed person, the inner body surface, and the parametric body model. Parametric Model-Conditioned Implicit Representation (PaMIR) [42] combines a parametric body model with a free-form deep implicit function to enhance generalization through regularization. Zhu et al. [205] utilized a ’project-predict-deform’ strategy that refines the SMPL model generated by human recovery methods using supervision from joints, silhouettes, and shading information. ICON [206] generates a stable, coarse human mesh using the SMPL model, then renders the front and back body normals, which provide rich texture details and combine with the original image. Finally, the body’s normal clothed normal, and Signed Distance Function (SDF) information are fed into an implicit MLP to obtain the final result. Subsequently, Xiu et al. proposed ECON [207], which integrates normal estimation, normal integration, and shape completion by using SMPL-X depth as a soft geometric constraint within the optimization equation of Normal Integration. It endeavors to maintain coherence with nearby surfaces during the integration of normals. Feng et al. [10] presented Disentangled Avatars (DELTA), a model representing humans with hybrid explicit-implicit 3D representations. GETAvatar [208] directly produces explicit textured 3D human meshes. GETAvatar creates an articulated 3D human representation with explicit surface modeling, enriches it with realistic surface details derived from 2D normal maps of 3D scan data, and utilizes a rasterization-based renderer for surface rendering. Diffusion Inpainting of Neural Avatars (DINAR) [209] combines neural textures with the SMPL-X body model and employs a latent diffusion model to recover textures of both seen and unseen regions, subsequently integrating these onto the base SMPL-X mesh. TransHuman [210] employs transformers to project the SMPL model into a canonical space and associates each output token with a deformable radiance field. This field encodes the query point in the observation space and is further utilized to integrate fine-grained information from reference images. Gaussian Splatting [220] is a technique for processing and rendering point clouds, using a Gaussian function to create an influence range for each point, resulting in a smoother and more natural representation in two-dimensional images, as shown in Fig.8(b). It has achieved good results in SLAM (Simultaneous Localization and Mapping) [221], generative human modeling [222], dynamic scene reconstruction [223], and multimodal generation [224]. Reconstruction based on Gaussian Splatting requires an initial point cloud input, which poses a challenge for human mesh recovery from image inputs. However, explicit models can provide initial points, serving as a stepping stone to make Gaussian splatting human mesh recovery from RGB inputs feasible. Animatable 3D Gaussian [211] learns human avatars from input images and poses by extending 3D Gaussian to dynamic human scenes. In the proposed framework, they model a set of skinned 3D Gaussian and a corresponding skeleton in canonical space and deform 3D Gaussian to posed space according to the input poses.